An electic current of 0.5 ampere is passed through acidulated water for 30 minutes.calculate the volume of hydrogen produced at NTP.(z for hydrogen=0.00001)
 

Grade:12

1 Answers

Aarti Gupta
9 years ago
Given – Current , I = 0.5A
Time ,t = 30min
Z for H2 = 0.00001
Volume ,V for hydrogen = ?

SOLUTION -
We know that ,
m = z *I * t
Putting the respective values,
m = 0.00001 * 0.5 * 30
m = 1.5 * 10-4 kg = 0.15g
now, density = mass / volume
therefore,
Volume,V = mass / density
Density of H2 gas at NTP = 0.089gm/l
So,
V = 0.15 / 0.089
V = 1.685ml or,
V of H2 produces at NTP = 1.66ml